\vskip1cm
\begin{thebibliography}{dlPG99}


\bibitem{An} Y. Andr\'e, Motifs de dimension finie (d'apr\`es S.-I. Kimura, P. O'Sullivan,...), S\'eminaire Bourbaki 2003/2004, Ast\'erisque 299 Exp. No. 929, viii, 115---145,

%\bibitem{And} Y. Andr\'e, Pour une th\'eorie inconditionnelle des motifs, Publications math\'ematiques de l'IHES 83 no 1 (1996), 5---49,

%\bibitem{A} D. Arapura, Motivation for Hodge cycles, Advances in Math. vol. 207 (2006), 762---781,

\bibitem{Beau0} A. Beauville, Some remarks on K\"ahler manifolds with $c_1=0$, in: Classification of algebraic and analytic manifolds (Katata, 1982), Birkh\"auser Boston, Boston 1983,

\bibitem{Beau1} A. Beauville, Vari\'et\'es K\"ahleriennes dont la premi\`ere classe de Chern est nulle, J. Differential Geom. 18 no. 4 (1983), 755---782,

\bibitem{Beau} A. Beauville, Sur l'anneau de Chow d'une vari\'et\'e ab\'elienne, Math. Ann. 273 (1986), 647---651,

\bibitem{Beau3} A. Beauville, On the splitting of the Bloch--Beilinson filtration, in: Algebraic cycles and motives (J. Nagel and C. Peters, editors), London Math. Soc. Lecture Notes 344, Cambridge 
University Press 2007,

\bibitem{Beau2} A. Beauville, Some surfaces with maximal Picard number, Journal de l'Ecole Polytechnique Tome 1 (2014), 101---116,

%\bibitem{BV} A. Beauville and C. Voisin, On the Chow ring of a $K3$ surface, J. Alg. Geom. 13 (2004), 417---426,

%\bibitem{Bl2} S. Bloch, Some elementary theorems about algebraic cycles on abelian varieties, Invent. Math. 37 (1976), 215---228,

\bibitem{B} S. Bloch, Lectures on algebraic cycles (2nd edition), Cambridge University Press New Mathematical Monographs 16, Cambridge 2010, 

%\bibitem{BKL} S. Bloch, A. Kas and D. Lieberman, Zero cycles on surfaces with $p_g=0$, Comp. Math. 33 no. 2 (1976), 135---145,

%\bibitem{BO} S. Bloch and A. Ogus, Gersten's conjecture and the homology of schemes, Ann. Sci. Ecole Norm. Sup. 4 (1974), 181---202,

\bibitem{BS} S. Bloch and V. Srinivas, Remarks on correspondences and algebraic cycles, American Journal of Mathematics Vol. 105, No 5 (1983), 1235---1253,

%\bibitem{BNS} S. Boissi\`ere, M. Nieper-Wi{\ss}kirchen and A. Sarti, Higher dimensional Enriques varieties and automorphisms of generalized Kummer varieties, Journal de Math\'ematiques Pures et Appliqu\'ees 95 (2011), 553---563,

%\bibitem{Bo} S. Boissi\`ere, Automorphismes naturels de l'espace de Douady de points sur une surface, Canad. J. Math. (2009),

%\bibitem{BCNS} S. Boissi\`ere, A. Cattaneo, M. Nieper-Wi{\ss}kirchen and A. Sarti, The automorphism group of the Hilbert scheme of two points on a generic projective $K3$ surface, 
%arXiv:1410.8387v2, in: K3 Surfaces and Their Moduli, Proceedings of the Schiermonnikoog conference 2014 (C. Faber, G. Farkas, G. van der Geer, editors), Progress in Maths 315,  Birkh\"auser 2016,

%\bibitem{BoSa} S. Boissi\`ere and A. Sarti, A note on automorphisms and birational transformations of holomorphic symplectic manifolds, Proc. Amer. Math. Soc. 140 (2012), 4053---4062,

%\bibitem{CM} 
%M. de Cataldo and L. Migliorini, The Chow groups and the motive of the Hilbert scheme of points on a
%surface, Journal of Algebra 251 no. 2 (2002), 824---848,

%\bibitem{ChMa} F. Charles and E. Markman, The Standard Conjectures for holomorphic symplectic varieties deformation equivalent to Hilbert schemes of $K3$ surfaces,
%Comp. Math. 149 no. 3 (2013), 481---494,

%\bibitem{D} P. Deligne, La conjecture de Weil pour les surfaces $K3$, Invent. Math. 15 (1972), 206---226,

%\bibitem{DM} C. Deninger and J. Murre, Motivic decomposition of abelian schemes and the Fourier transform. J. reine u.
%angew. Math. 422 (1991), 201---219,

%\bibitem{EPW} D. Eisenbud, S. Popescu and C. Walter, Lagrangian subbundles and codimension $3$ subcanonical subschemes, Duke Math. J. 107(3) (2001), 427---467,

%\bibitem{LFu} L. Fu, Decomposition of small diagonals and Chow rings of hypersurfaces and Calabi--Yau complete intersections, Advances in Mathematics (2013), 894---924,

%\bibitem{LFu} L. Fu, Beauville-Voisin conjecture for generalized Kummer varieties, International Mathematics Research Notices 12 (2015), 3878---3898,
%2014,

\bibitem{LFu3} L. Fu, Classification of polarized symplectic automorphisms of Fano varieties of cubic fourfolds, Glasgow Math. Journal 58 No. 1 (2016), 17---37,


%\bibitem{LFu2} L. Fu, On the action of symplectic automorphisms on the $CH_0$--groups of some hyper-K\"ahler fourfolds, Math. Z. 280 (2015), 307---334,

\bibitem{FTV} L. Fu, Z. Tian and C. Vial, Motivic hyperk\"ahler resolution conjecture for generalized Kummer varieties, arXiv:1608.04968,

%\bibitem{F} W. Fulton, Intersection theory, Springer--Verlag Ergebnisse der Mathematik, Berlin Heidelberg New York Tokyo 1984,


%\bibitem{HI} A. Hirschowitz and J. Iyer, Hilbert schemes of fat r--planes and the triviality of Chow groups of complete intersections. In: Vector
%bundles and complex geometry, Contemp. Math. 522, Amer. Math. Soc., Providence (2010),

%\bibitem{Hu} D. Huybrechts, Symplectic automorphisms of $K3$ surfaces of arbitrary order, Math. Res. Letters 19 (2012), 947---951,

%\bibitem{Hui} M. Huibregtse, Translation by torsion points and rational equivalence of $0$--cycles on abelian varieties, Rocky Mountain J. Math. 18 no. 1 (1988), 147---155,

%\bibitem{WHu} W. Hu, The Chow group of zero cycles for the quotient of certain Calabi--Yau varieties, Houston J. of Math. Vol 38, No 1 (2012), 55---67,


\bibitem{Iv} F. Ivorra, Finite dimensional motives and applications (following S.-I. Kimura, P. O'Sullivan and others), in:
        Autour des motifs, Asian-French summer school on algebraic geometry and number theory,
       Volume III, Panoramas et synth\`eses, Soci\'et\'e math\'ematique de France 2011,

% \bibitem{Iy} J. Iyer, Murre's conjectures and explicit Chow--K\"unneth projectors for varieties with a nef tangent bundle, Transactions of the Amer. Math. Soc. 361 (2008), 1667---1681,

%\bibitem{Iy2} J. Iyer, Absolute Chow--K\"unneth decomposition for rational homogeneous bundles and for log homogeneous varieties, Michigan Math. Journal
% Vol.60, 1 (2011), 79---91,
          
%\bibitem{J} U. Jannsen, Mixed motives and algebraic K--theory, Springer Lecture Notes in Mathematics 1400 (1990),

%\bibitem{J} U. Jannsen, 
%Motives, numerical equivalence, and semi-simplicity, Invent. Math. 107(3) (1992), 447---452, 

\bibitem{J2} U. Jannsen, Motivic sheaves and filtrations on Chow groups, in: Motives (U. Jannsen et alii, eds.), Proceedings of Symposia in Pure Mathematics Vol. 55 (1994), Part 1,  

%\bibitem{J3} U. Jannsen, Equivalence relations on algebraic cycles, in: The arithmetic and geometry
%of algebraic cycles (B. Gordon et alii, eds.), Banff Conference 1998, Kluwer,

\bibitem{J4} U. Jannsen, On finite--dimensional motives and Murre's conjecture, in: Algebraic cycles and motives (J. Nagel and C. Peters, editors), Cambridge University Press, Cambridge 2007,

\bibitem{KMP} B. Kahn, J. Murre and C. Pedrini, On the transcendental part of the motive of a surface, in: Algebraic cycles and motives (J. Nagel and C. Peters, editors), Cambridge University Press, Cambridge 2007,

%\bibitem{KS} B. Kahn and R. Sebastian, Smash--nilpotent cycles on abelian 3--folds, Math. Res. Letters 16 (2009), 1007---1010,

\bibitem{KS} T. Katsura and T. Shioda, On Fermat varieties, Tohoku Math. J. Vol. 31 No. 1 (1979), 97---115,

\bibitem{Kawa} K. Kawatani, On the birational geometry for irreducible symplectic $4$--folds related to the Fano schemes of lines, arXiv:0906.0654,

\bibitem{Kim} S. Kimura, Chow groups are finite dimensional, in some sense,
Math. Ann. 331 (2005), 173---201,

%\bibitem{KV} S. Kimura and A. Vistoli, Chow rings of infinite symmetric products, Duke Math. J. 85 (1996), 411---430,

%\bibitem{K0} S. Kleiman, Algebraic cycles and the Weil conjectures, in: Dix expos\'es sur la cohomologie des sch\'emas, North--Holland Amsterdam, 1968, 359---386, 

%\bibitem{K} S. Kleiman, The standard conjectures, in: Motives (U. Jannsen et alii, eds.), Proceedings of Symposia in Pure Mathematics Vol. 55 (1994), Part 1, 

%\bibitem{Ko} K. Koike, Elliptic $K3$ surfaces admitting a Shioda--Inose structure, Comment. Math. Univ. St. Pauli 61 No 1 (2012), 77---86,


%\bibitem{small} R. Laterveer, Algebraic varieties with small Chow groups, J. Math. Kyoto Univ. Vol. 38 No 4 (1998), 673---694,

%\bibitem{yet} R. Laterveer, Yet another version of Mumford's theorem, Archiv Math. (2015),   

%\bibitem{moi2} R. Laterveer, A brief note concerning hard Lefschetz for Chow groups, to appear in Canadian Math. Bulletin, 

\bibitem{moi} R. Laterveer, Some results on a conjecture of Voisin for surfaces of geometric genus one, Boll. Unione Mat. Italiana 9 no. 4 (2016), 435---452, 

%\bibitem{moimum} R. Laterveer, On a multiplicative version of Mumford's theorem, Abh. Math. Sem. Univ. Hambg 86(1) (2016), 89---96,

%\bibitem{moimult} R. Laterveer, On a multiplicative version of Bloch's conjecture, to appear in Beitr\"age zum Algebra und Geometrie,

%\bibitem{moi} R. Laterveer, Bloch's conjecture for generalized Kummer varieties, submitted,

\bibitem{tod} R. Laterveer, Algebraic cycles and Todorov surfaces, to appear in Kyoto Journal of Mathematics, arXiv:1609.09629,

\bibitem{fano} R. Laterveer, A remark on the motive of the Fano variety of lines on a cubic, Ann. Math. Qu\'ebec 41 no. 1 (2017), 141---154,

%\bibitem{cubic4} R. Laterveer, A family of cubic fourfolds with finite--dimensional motive, submitted,

%\bibitem{EPW} R. Laterveer, Algebraic cycles on a very special EPW sextic, submitted,

%\bibitem{BlochHK4} R. Laterveer, Bloch's conjecture for certain hyperk\"ahler fourfolds, and EPW sextics, submitted,

%\bibitem{moi} R. Laterveer, A family of $K3$ surfaces having finite--dimensional motive, to appear in Archiv der Math.,

\bibitem{des} R. Laterveer, Some desultory remarks concerning algebraic cycles and Calabi--Yau threefolds, Rend. Circ. Mat. Palermo 65 (2) (2016), 
333---344,

%\bibitem{exCY} R. Laterveer, A family of Calabi--Yau threefolds with finite--dimensional motive, submitted to Tokyo Math. J.,

%\bibitem{Lew} J. Lewis, Cylinder homomorphisms and Chow groups, Math. Nachr. 160 (1993), 205---221,

%\bibitem{Li} C. Liedtke, Supersingular $K3$ surfaces are unirational, Invent. Math. 200 (2015), 979---1014,

%\bibitem{Lin} H. Lin, On the Chow group of zero--cycles of a generalized Kummer variety, Advances in Mathematics 298 (2016), 448---472,

%\bibitem{Ma} S. Ma, On $K3$ surfaces which dominate Kummer surfaces, Proceedings of the American Math. Soc. 141 No 1 (2013), 131---137,

%\bibitem{MTW} G. Mongardi, K. Tari and M. Wandel, Automorphisms of generalized Kummer fourfolds, arXiv:1512.00225v2,

%\bibitem{Mum} D. Mumford, Rational equivalence of $0$--cycles on surfaces, J. Math. Kyoto Univ. Vol. 9 No 2 (1969), 195---204,

\bibitem{Mur} J. Murre, On a conjectural filtration on the Chow groups of an algebraic variety, parts I and II, Indag. Math. 4 (1993), 177---201,

\bibitem{MNP} J. Murre, J. Nagel and C. Peters, Lectures on the theory of pure motives, Amer. Math. Soc. University Lecture Series 61, Providence 2013,

%\bibitem{Murt} V. Murty, Algebraic cycles on abelian varieties, Duke Math. J. 50 (1983), 487---504,

\bibitem{Nam} Y. Namikawa, Deformation theory of singular symplectic $n$--folds,
Math. Ann. 319 (2001), 597---623,

%\bibitem{OG} K. O'Grady, Decomposable cycles and Noether--Lefschetz loci, arXiv:1502.07897v2,

%\bibitem{OG2} K. O'Grady, Involutions and linear systems on holomorphic symplectic manifolds, Geom. Funct. Anal. 15 no 6 (2005), 1223---1274,

%\bibitem{OG} K. O'Grady, Irreducible symplectic $4$--folds and Eisenbud--Popescu--Walter sextics, Duke Math. J. 134(1) (2006), 99---137,

%\bibitem{OG4} K. O'Grady, EPW--sextics: taxonomy, Manuscripta Math. 138 1 (2012), 221---272,

%\bibitem{OG3} K. O'Grady, Double covers of EPW--sextics, Michigan Math. J. 62 (2013), 143---184,

%\bibitem{Ogu} K. Oguiso, A remark on dynamical degrees of automorphisms of hyperk\"ahler manifolds, Manuscripta Math. 130 (2009), 101---111,

%\bibitem{P} C. Pedrini, On the finite dimensionality of a $K3$ surface, Manuscripta Mathematica 138 (2012), 59---72,

%\bibitem{PW} C. Pedrini and C. Weibel, Some surfaces of general type for which Bloch's conjecture holds, to appear in: Period Domains, Algebraic Cycles, and Arithmetic, Cambridge Univ. Press, 2015,

\bibitem{Rie} U. Rie\ss, On the Chow ring of birational irreducible symplectic varieties, Manuscripta Math. 145 (2014), 473---501,

%\bibitem{Rito} C. Rito, A note on Todorov surfaces, Osaka Journal of Math. 46(3) (2009), 685---693,

%\bibitem{Ro} A.A. Rojtman, The torsion of the group of 0--cycles modulo rational equivalence, Annals of Mathematics 111 (1980), 553---569,

\bibitem{Sc} T. Scholl, Classical motives, in: Motives (U. Jannsen et alii, eds.), Proceedings of Symposia in Pure Mathematics Vol. 55 (1994), Part 1, 

%\bibitem{Seb} R. Sebastian, Smash nilpotent cycles on varieties dominated by products of curves, Comp. Math. 149 (2013), 1511---1518,

%\bibitem{Seb2} R. Sebastian, Examples of smash nilpotent cycles on rationally connected varieties, Journal of Algebra 438 (2015), 119---129,

\bibitem{SV} M. Shen and C. Vial, The Fourier transform for certain hyperK\"ahler fourfolds, Memoirs of the AMS 240 (2016), no.1139,

\bibitem{SV2} M. Shen and C. Vial, The motive of the Hilbert cube $X^{[3]}$, Forum Math Sigma 4 (2016), 55 pages, 

\bibitem{Shi} T. Shioda, The Hodge conjecture for Fermat varieties, Math. Ann. 245 (1979), 175---184,

%\bibitem{V2} C. Vial, Projectors on the intermediate algebraic Jacobians, New York J. Math. 19 (2013), 793---822,

%\bibitem{V3} C. Vial, Remarks on motives of abelian type, to appear in Tohoku Math. J.,

%\bibitem{V4} C. Vial, Niveau and coniveau filtrations on cohomology groups and Chow groups, Proceedings of the LMS 106(2) (2013), 410---444,

%\bibitem{V5} C. Vial, Chow--K\"unneth decomposition for $3$-- and $4$--folds fibred by varieties with trivial Chow group of zero--cycles, J. Alg. Geom. 24 (2015), 51---80,

\bibitem{V6} C. Vial, On the motive of some hyperk\"ahler varieties, J. f\"ur Reine u. Angew. Math. 725 (2017), 235---247,

%\bibitem{Voe} V. Voevodsky, A nilpotence theorem for cycles algebraically equivalent to zero, Internat. Math. Research Notices 4 (1995), 187---198,

\bibitem{V9} C. Voisin, Remarks on zero--cycles of self--products of varieties, in: Moduli of vector bundles, Proceedings of the Taniguchi Congress  (M. Maruyama,  ed.), Marcel Dekker New York Basel Hong Kong 1994,

%\bibitem{V10} C. Voisin, Remarks on filtrations on Chow groups and the Bloch conjecture,
%Annali di matematica pura ed applicata 183 (2004), 421---438,

%\bibitem{V20} C. Voisin, Miroirs et involutions sur les surfaces K3, Ast\'erisque 218 (1993), 273---323,

%\bibitem{V12} C. Voisin, Sur les z\'ero--cycles de certaines hypersurfaces munies d'un automorphisme, Annali della Scuola Norm. Sup. di Pisa Vol. 29 (1993), 473---492,

%\bibitem{V12} C. Voisin, On the Chow ring of certain algebraic hyper--K\"ahler manifolds, Pure and Applied Mathematics Quarterly 4 no. 3 (2008), 613---649,

%\bibitem{V13} C. Voisin, Chow rings and decomposition theorems for $K3$ surfaces and Calabi--Yau hypersurfaces, Geom. Topol.
%16 (2012), 433---473,

%\bibitem{V11} C. Voisin, Symplectic involutions of $K3$ surfaces act trivially on $CH_0$, Documenta Math. 17 (2012), 851---860,

%\bibitem{V0} C. Voisin, The generalized Hodge and Bloch conjectures are equivalent for general complete intersections, Ann. Sci. Ecole Norm. Sup. 46, fascicule 3 (2013), 449---475,

%\bibitem{V1} C. Voisin, The generalized Hodge and Bloch conjectures are equivalent for general complete intersections, II, J. Math. Sci. Univ. Tokyo  22 (2015), 491---517,

%\bibitem{V8} C. Voisin, Bloch's conjecture for Catanese and Barlow surfaces, J. Differential Geometry 97 (2014), 149---175,

\bibitem{Vo} C. Voisin, Chow Rings, Decomposition of the Diagonal, and the Topology of Families, Princeton University Press, Princeton and Oxford, 2014,

\bibitem{V14} C. Voisin, Remarks and questions on coisotropic subvarieties and $0$--cycles of hyper--K\"ahler varieties, in: K3 Surfaces and Their Moduli, Proceedings of the Schiermonnikoog conference 2014 (C. Faber, G. Farkas, G. van der Geer, editors), Progress in Maths 315,  Birkh\"auser 2016.

\end{thebibliography}

\documentclass[a4paper,12pt,final]{amsart}
\usepackage{times,a4wide,mathrsfs,amssymb,amsmath,amsthm}

\newcommand{\C}{\mathbb{C}}

\newcommand{\LLL}{\mathbb{L}}
\newcommand{\QQ}{\mathbb{Q}}
\newcommand{\NN}{\mathbb{N}}
\newcommand{\PP}{\mathbb{P}}

\newcommand{\MM}{\mathcal M}

\newcommand{\gr}{\hbox{Gr}}

\newcommand{\rom}{\romannumeral}

\DeclareMathOperator{\aut}{Aut}

\newtheorem{theorem}{Theorem}[section]

\newtheorem{corollary}[theorem]{Corollary}
\newtheorem{proposition}[theorem]{Proposition}
\newtheorem{conjecture}[theorem]{Conjecture}
\newtheorem{remark}[theorem]{Remark}
\newtheorem{definition}[theorem]{Definition}
\newtheorem{convention}{Conventions}

\newtheorem{nonumbering}{Theorem}

\newtheorem{nonumberingc}{Corollary}

\newtheorem{nonumberingt}{Acknowledgements}

\begin{document}
\author[Robert Laterveer]
{Robert Laterveer}

\address{Institut de Recherche Math\'ematique Avanc\'ee,
CNRS -- Universit\'e 
de Strasbourg,\
7 Rue Ren\'e Des\-car\-tes, 67084 Strasbourg CEDEX,
FRANCE.}
\email{robert.laterveer@math.unistra.fr}

\title{On Voisin's conjecture for zero--cycles on hyperk\"ahler varieties}

\begin{abstract} Motivated by the Bloch--Beilinson conjectures, Voisin has made a conjecture concerning zero--cycles on self--products of Calabi--Yau varieties. We reformulate Voisin's conjecture in the setting of hyperk\"ahler varieties, and we prove this reformulated conjecture for one family of hyperk\"ahler fourfolds.
\end{abstract}

\keywords{Algebraic cycles, Chow groups, motives, Bloch's conjecture, Bloch--Beilinson filtration, hyperk\"ahler varieties, multiplicative Chow--K\"unneth decomposition, splitting property, finite--dimensional motive}
\subjclass[2010]{Primary 14C15, 14C25, 14C30.}

\maketitle

\section{Introduction}

Let $X$ be a smooth projective variety over $\C$, and let $A^i(X):=CH^i(X)_{\QQ}$ denote the Chow groups of $X$ (i.e. the groups of codimension $i$ algebraic cycles on $X$ with $\QQ$--coefficients, modulo rational equivalence). As is well--known, the field of algebraic cycles is rife with open questions \cite{B}, \cite{J2}, \cite{MNP}, \cite{Vo}. To wit, there is the following conjecture formulated by Voisin, which can be seen as a version of Bloch's conjecture for varieties of geometric genus one:

\begin{conjecture}[Voisin \cite{V9}]\label{conjvois} Let $X$ be a smooth projective complex variety of dimension $n$ with $h^{j,0}(X)=0$ for $0<j<n$ and $p_g(X)=1$.  
For any zero--cycles $a,a^\prime\in A^n(X)$ of degree zero, we have
  \[ a\times a^\prime=(-1)^n a^\prime\times a\ \ \ \hbox{in}\ A^{2n}(X\times X)\ .\]
  (Here $a\times a^\prime$ is short--hand for the cycle class $(p_1)^\ast(a)\cdot(p_2)^\ast(a^\prime)\in A^{2n}(X\times X)$, where $p_1, p_2$ denote projection on the first, resp. second factor.)
   \end{conjecture}
   
 Conjecture \ref{conjvois} is still wide open for a general $K3$ surface (on the positive side, cf. \cite{V9}, \cite{moi}, \cite{des}, \cite{tod} for some cases where this conjecture is verified).

Let us now suppose that $X$ is a hyperk\"ahler variety (i.e., a projective irreducible holomorphic symplectic manifold, cf. \cite{Beau0}, \cite{Beau1}). Conjecture \ref{conjvois} does not apply verbatim to $X$ (the Calabi--Yau condition 
$h^{j,0}(X)=0$ for $0<j<n$ is not satisfied), yet one can adapt conjecture \ref{conjvois} to make sense for $X$. For this adaptation, we will optimistically assume the Chow ring of $X$ has a bigraded ring structure $A^\ast_{(\ast)}(X)$, where each $A^i(X)$ splits into pieces
  \[ A^i(X) =\bigoplus_j A^i_{(j)}(X)\ .\]
 (Conjecturally, such a splitting exists for all hyperk\"ahler varieties, and the piece $A^i_{(j)}(X)$ should be isomorphic to the graded $\gr^j_F A^i(X)$ for the conjectural Bloch--Beilinson filtration \cite{Beau3}. Using the concept of ``(weak) multiplicative Chow--K\"unneth decompositions'', a bigraded ring structure has been constructed for Hilbert schemes of length $n$ subschemes of $K3$ surfaces \cite{V6}, for generalized Kummer varieties \cite{FTV}, and for Fano varieties of lines in very general cubic fourfolds \cite{SV}. For a different approach to the construction of a bigraded ring structure on the Chow ring of hyperk\"ahler varieties, cf. \cite{V14}.) 
 
 Since the piece $A^i_{(j)}(X)$ should be related to the cohomology group $H^{2i-j}(X)$, and 
   \[ \wedge^2  H^{s}(X)\ \subset\ H^{2s}(X\times X) \]
   should be supported on a divisor for any $s$ (in view of the generalized Hodge conjecture), we arrive at the following version of conjecture \ref{conjvois}:

\begin{conjecture}\label{conjHK} Let $X$ be a hyperk\"ahler variety of dimension $2m$. Let $a,a^\prime\in A^{2m}_{(j)}(X)$. Then
  \[ a\times a^\prime -a^\prime\times a=0\ \ \ \hbox{in}\ A^{4m}_{}(X\times X)\ .\]
  \end{conjecture}
  
  (We note that in conjecture \ref{conjHK}, we silently presuppose that $A^{2m}_{(j)}(X)=0$ for $j$ odd. This is expected to hold for any hyperk\"ahler variety, and is known unconditionally for the above--mentioned cases where a bigraded ring structure exists.)

  The main result of this note is that conjecture \ref{conjHK} is true for a certain family of hyperk\"ahler fourfolds:
  
  \begin{nonumbering}[=theorem \ref{main}] Let $Y\subset\PP^5(\C)$ be a smooth cubic fourfold defined by an equation
    \[ f(X_0,X_1,X_2)+ g(X_3,X_4,X_5)=0\ ,\]
    where $f$ and $g$ define smooth plane curves. Let $X=F(Y)$ be the Fano variety of lines in $Y$. Then for any $a,a^\prime\in A^4_{(j)}(X)$, there is equality
     \[ a\times a^\prime -a^\prime\times a=0\ \ \ \hbox{in}\ A^{8}_{}(X\times X)\ .\]
    \end{nonumbering} 
    
    Here the notation $A^\ast_{(\ast)}(X)$ refers to the Fourier decomposition of the Chow ring constructed by Shen--Vial \cite{SV}. (We mention in passing that for $X$ as in theorem \ref{main}, it is not yet known whether the pieces $A^\ast_{(\ast)}(X)$ fit together to form a bigraded ring, cf. remark \ref{pity}.) 
    
    The proof of theorem \ref{main} relies on the theory of finite--dimensional motives \cite{Kim}.
    
   As a corollary to theorem \ref{main}, we find that a certain instance of the generalized Hodge conjecture is verified on $X\times X$: 
   
   \begin{nonumberingc}[=corollary \ref{cor}] Let $X$ be as in theorem \ref{main}. The Hodge sub--structure
     \[ \wedge^2 H^4(X,\QQ)\ \subset\ H^8(X\times X,\QQ) \]
     is supported on a divisor.
   \end{nonumberingc}

 \vskip0.6cm

\begin{convention} In this article, the word {\sl variety\/} will refer to a reduced irreducible scheme of finite type over $\C$. A {\sl subvariety\/} is a (possibly reducible) reduced subscheme which is equidimensional. 

{\bf All Chow groups will be with rational coefficients}: we will denote by $A_j(X)$ the Chow group of $j$--dimensional cycles on $X$ with $\QQ$--coefficients; for $X$ smooth of dimension $n$ the notations $A_j(X)$ and $A^{n-j}(X)$ are used interchangeably. 

The notations $A^j_{hom}(X)$, $A^j_{AJ}(X)$ will be used to indicate the subgroups of homologically trivial, resp. Abel--Jacobi trivial cycles.
For a morphism $f\colon X\to Y$, we will write $\Gamma_f\in A_\ast(X\times Y)$ for the graph of $f$.
The contravariant category of Chow motives (i.e., pure motives with respect to rational equivalence as in \cite{Sc}, \cite{MNP}) will be denoted $\MM_{\rm rat}$.



We will write $H^j(X)$ 
to indicate singular cohomology $H^j(X,\QQ)$.

\end{convention}

\section{Preliminaries}

\subsection{Finite--dimensional motives}

We refer to \cite{Kim}, \cite{An}, \cite{Iv}, \cite{J4}, \cite{MNP} for the definition of finite--dimensional motive. 
An essential property of varieties with finite--dimensional motive is embodied by the nilpotence theorem:

\begin{theorem}[Kimura \cite{Kim}]\label{nilp} Let $X$ be a smooth projective variety of dimension $n$ with finite--dimensional motive. Let $\Gamma\in A^n(X\times X)_{}$ be a correspondence which is numerically trivial. Then there is $N\in\NN$ such that
     \[ \Gamma^{\circ N}=0\ \ \ \ \in A^n(X\times X)_{}\ .\]
\end{theorem}

 Actually, the nilpotence property (for all powers of $X$) could serve as an alternative definition of finite--dimensional motive, as shown by Jannsen \cite[Corollary 3.9]{J4}.
Conjecturally, any variety has finite--dimensional motive \cite{Kim}. We are still far from knowing this, but at least there are quite a few non--trivial examples.

\subsection{MCK decomposition}

\begin{definition}[Murre \cite{Mur}] Let $X$ be a projective quotient variety of dimension $n$. We say that $X$ has a {\em CK decomposition\/} if there exists a decomposition of the diagonal
   \[ \Delta_X= \pi_0+ \pi_1+\cdots +\pi_{2n}\ \ \ \hbox{in}\ A^n(X\times X)\ ,\]
  such that the $\pi_i$ are mutually orthogonal idempotents and $(\pi_i)_\ast H^\ast(X)= H^i(X)$.
  
  (NB: ``CK decomposition'' is shorthand for ``Chow--K\"unneth decomposition''.)
\end{definition}

\begin{remark}
\normalfont The existence of a CK decomposition for any smooth projective variety is part of Murre's conjectures \cite{Mur}, \cite{J2}. 
\end{remark}

\begin{definition}[Shen--Vial \cite{SV}]
\normalfont Let $X$ be a projective quotient variety of dimension $n$. Let $\Delta_X^{sm}\in A^{2n}(X\times X\times X)$ be the class of the small diagonal
  \[ \Delta_X^{sm}:=\bigl\{ (x,x,x)\ \vert\ x\in X\bigr\}\ \subset\ X\times X\times X\ .\]
  An {\em MCK decomposition\/} is a CK decomposition $\{\pi^X_i\}$ of $X$ that is {\em multiplicative\/}, i.e. it satisfies
  \[ \pi^X_k\circ \Delta_X^{sm}\circ (\pi^X_i\times \pi^X_j)=0\ \ \ \hbox{in}\ A^{2n}(X\times X\times X)\ \ \ \hbox{for\ all\ }i+j\not=k\ .\]
  
 (NB: ``MCK decomposition'' is shorthand for ``multiplicative Chow--K\"unneth decomposition''.) 
  
 A {\em weak MCK decomposition\/} is a CK decomposition $\{\pi^X_i\}$ of $X$ that satisfies
    \[ \Bigl(\pi^X_k\circ \Delta_X^{sm}\circ (\pi^X_i\times \pi^X_j)\Bigr){}_\ast (a\times b)=0 \ \ \ \hbox{for\ all\ } a,b\in\ A^\ast(X)\ .\]
  \end{definition}
  
  \begin{remark}
  \normalfont The small diagonal (seen as a correspondence from $X\times X$ to $X$) induces the {\em multiplication morphism\/}
    \[ \Delta_X^{sm}\colon\ \  h(X)\otimes h(X)\ \to\ h(X)\ \ \ \hbox{in}\ \MM_{\rm rat}\ .\]
 Suppose $X$ has a CK decomposition
  \[ h(X)=\bigoplus_{i=0}^{2n} h^i(X)\ \ \ \hbox{in}\ \MM_{\rm rat}\ .\]
  By definition, this decomposition is multiplicative if for any $i,j$ the composition
  \[ h^i(X)\otimes h^j(X)\ \to\ h(X)\otimes h(X)\ \xrightarrow{\Delta_X^{sm}}\ h(X)\ \ \ \hbox{in}\ \MM_{\rm rat}\]
  factors through $h^{i+j}(X)$.
  
  If $X$ has a weak MCK decomposition, then setting
    \[ A^i_{(j)}(X):= (\pi^X_{2i-j})_\ast A^i(X) \ ,\]
    one obtains a bigraded ring structure on the Chow ring: that is, the intersection product sends $A^i_{(j)}(X)\otimes A^{i^\prime}_{(j^\prime)}(X) $ to  $A^{i+i^\prime}_{(j+j^\prime)}(X)$.
    
      It is expected (but not proven !) that for any $X$ with a weak MCK decomposition, one has
    \[ A^i_{(j)}(X)\stackrel{??}{=}0\ \ \ \hbox{for}\ j<0\ ,\ \ \ A^i_{(0)}(X)\cap A^i_{hom}(X)\stackrel{??}{=}0\ ;\]
    this is related to Murre's conjectures B and D, that have been formulated for any CK decomposition \cite{Mur}.

  The property of having an MCK decomposition is severely restrictive, and is closely related to Beauville's ``(weak) splitting property'' \cite{Beau3}. For more ample discussion, and examples of varieties with an MCK decomposition, we refer to \cite[Section 8]{SV}, as well as \cite{V6}, \cite{SV2}, \cite{FTV}.
    \end{remark}

In what follows, we will make use of the following: 

\begin{theorem}[Shen--Vial \cite{SV}]\label{fanomck} Let $Y\subset\PP^5(\C)$ be a smooth cubic fourfold, and let $X:=F(Y)$ be the Fano variety of lines in $Y$. There exists a CK decomposition $\{\pi^X_i\}$ for $X$, and 
  \[ (\pi^X_{2i-j})_\ast A^i(X) = A^i_{(j)}(X)\ ,\]
  where the right--hand side denotes the splitting of the Chow groups defined in terms of the Fourier transform as in \cite[Theorem 2]{SV}. Moreover, we have
  \[ A^i_{(j)}(X)=0\ \ \ \hbox{for\ }j<0\ \hbox{and\ for\ }j>i\ .\]
  
  In case $Y$ is very general, the Fourier decomposition $A^\ast_{(\ast)}(X)$ forms a bigraded ring, and hence
  $\{\pi^X_i\}$ is a weak MCK decomposition.
    \end{theorem}

\begin{proof} (A remark on notation: what we denote $A^i_{(j)}(X)$ is denoted $CH^i(X)_j$ in \cite{SV}.)

The existence of a CK decomposition $\{\pi^X_i\}$ is \cite[Theorem 3.3]{SV}, combined with the results in \cite[Section 3]{SV} to ensure that the hypotheses of \cite[Theorem 3.3]{SV} are satisfied. According to \cite[Theorem 3.3]{SV}, the given CK decomposition agrees with the Fourier decomposition of the Chow groups. The ``moreover'' part is because the $\{\pi^X_i\}$ are shown to satisfy Murre's conjecture B \cite[Theorem 3.3]{SV}.

The statement for very general cubics is \cite[Theorem 3]{SV}.
    \end{proof}

\begin{remark}\label{pity}
\normalfont Unfortunately, it is not yet known that the Fourier decomposition of \cite{SV} induces a bigraded ring structure on the Chow ring for {\em all\/} Fano varieties of smooth cubic fourfolds. For one thing, it has not yet been proven that $A^2_{(0)}(X)\cdot A^2_{(0)}(X)\subset A^4_{(0)}(X)$ (cf. \cite[Section 22.3]{SV} for discussion).
\end{remark}



\subsection{Fano varieties of cubic fourfolds}

Let $X$ be the Fano variety of lines on a smooth cubic fourfold. As we have seen (theorem \ref{fanomck}), the Chow ring of $X$ splits into pieces $A^i_{(j)}(X)$.
The magnum opus \cite{SV} contains a detailed analysis of the multiplicative behaviour of these pieces. Here are the relevant results we will be needing:

\begin{theorem}[Shen--Vial \cite{SV}]\label{chowringfano} Let $Y\subset\PP^5(\C)$ be a smooth cubic fourfold, and let $X:=F(Y)$ be the Fano variety of lines in $Y$. 

\noindent
(\rom1) There exists $\ell\in A^2_{(0)}(X)$ such that intersecting with $\ell$ induces an isomorphism
  \[ \cdot\ell\colon\ \ \ A^2_{(2)}(X)\ \xrightarrow{\cong}\ A^4_{(2)}(X)\ .\]

\noindent
(\rom2) Intersection product induces a surjection
  \[ A^2_{(2)}(X)\otimes A^2_{(2)}(X)\ \twoheadrightarrow\ A^4_{(4)}(X)\ .\]
\end{theorem} 
     
 \begin{proof} Statement (\rom1) is \cite[Theorem 4]{SV}. Statement (\rom2) is \cite[Proposition 20.3]{SV}.
  \end{proof}

\section{Main result}

\begin{theorem}\label{main}  Let $Y\subset\PP^5(\C)$ be a smooth cubic fourfold defined by an equation
    \[ f(X_0,X_1,X_2)+ g(X_3,X_4,X_5)=0\ ,\]
    where $f$ and $g$ define smooth plane curves. Let $X=F(Y)$ be the Fano variety of lines in $Y$. Then for any $a,a^\prime\in A^4_{(j)}(X)$, there is equality
     \[ a\times a^\prime -a^\prime\times a=0\ \ \ \hbox{in}\ A^{8}_{(2j)}(X\times X)\ .\]   
  \end{theorem}     
 
 \begin{proof}    
  Let $u_Y\colon Y\to Y$ be the automorphism defined by the automorphism of $\PP^5(\C)$
    \[  [X_0\colon\cdots\colon X_5]\ \mapsto\ [X_0\colon X_1\colon X_2\colon \omega X_3\colon \omega X_4\colon \omega X_5]\ ,\]
    where $\omega$ is a primitive third root of unity. Let $u_X\in\aut(X)$ denote the automorphism induced by $u_Y$. As explained in \cite[Proof of Proposition 12]{Beau2}, $u_X$ is a symplectic automorphism, and the fixed locus $B\subset X$ of $u_X$ is isomorphic to $E\times E^\prime$, where $E$ and $E^\prime$ are the elliptic curves 
    \[ \begin{split} E&:= Y\cap \{ X_0=X_1=X_2=0\}\ ,\\
                       E^\prime&:= Y\cap \{ X_3=X_4=X_5=0\}\ .\\
                       \end{split}\]
  What's more, as $u_X$ is symplectic, $\tau\colon B\to X$ is the inclusion of a symplectic submanifold and so restriction induces an isomorphism
   \[    \tau^\ast    \colon\ \ H^{2,0}(X)\ \to\ H^{2,0}(B)\ .\]   
  This implies there is also an isomorphism of transcendental lattices
    \[    \tau^\ast    \colon\ \ H^{2}_{tr}(X)\ \to\ H^{2}_{tr}(B)\ \]
    (here $H^2_{tr}()\subset H^2()$ is defined as the smallest Hodge sub--structure containing $H^{2,0}$).   
   A bit more formally, this means there is an isomorphism of homological motives
   \begin{equation}\label{isomot} \Gamma= {}^t\Gamma_\tau+ \Gamma^\prime \colon\ \ \ (X,\pi^X_{2},0)\ \xrightarrow{\cong}\ (B,\pi^B_{2,tr},0) \oplus  \LLL^{\oplus \rho}\ \ \ \hbox{in}\ \MM_{\rm hom}\ .\end{equation}
  Here $\LLL$ denotes the Lefschetz motive, and $\rho$ is the dimension of the N\'eron--Severi group of $X$. The projector $\pi^X_2$ is as in theorem \ref{fanomck}, and $\pi^B_{2,tr}$ is a projector defining the transcendental part of the motive of a surface as in \cite{KMP}.
   
   As both sides of (\ref{isomot}) are finite--dimensional motives (for the left--hand side, this follows from proposition \ref{fanofindim}), we can upgrade the isomorphism (\ref{isomot}) to an isomorphism of Chow motives
   \begin{equation}\label{isomotch} \Gamma\colon\ \ \ (X,\pi^X_{2},0)\ \xrightarrow{\cong}\ (B,\pi^B_{2,tr},0)\oplus \LLL^{\oplus \rho}\ \ \ \hbox{in}\ \MM_{\rm rat}\ .\end{equation}  
               Since isomorphic Chow motives have isomorphic Chow groups, the isomorphism (\ref{isomotch}) implies there is an induced isomorphism
      \[ (\pi^B_{2,tr} \circ \Gamma \circ \pi^X_2)_\ast  \colon\ \ \ (\pi^X_2)_\ast A^2_{hom}(X)\ \xrightarrow{\cong}\  (\pi^B_{2,tr})_\ast A^2_{hom}(B)\oplus 
      A^2_{hom}(\LLL^{\oplus \rho})\ .\]
      But $A^2_{(2)}()\subset A^2_{hom}()$, and $A^2_{hom}(\LLL)=0$, and so we find that   
            restriction induces an isomorphism
   \[  (\pi^B_{2,tr}\circ {}^t \Gamma_\tau)_\ast\colon\ \ \ A^2_{(2)}(X)\ \xrightarrow{\cong}\  ( \pi^B_{2,tr})_\ast A^2_{hom}(B)=A^2_{(2)}(B)=A^2_{AJ}(B)\ .\]
   (Here, the notation $A^2_{(2)}(B)$ refers to Beauville's splitting of the Chow ring of abelian varieties \cite{Beau}.)
   
  We are now in position to state a result that will serve as an intermediate step towards proving theorem \ref{main}:
  
  \begin{proposition}\label{intermed} Let $X$ be as in theorem \ref{main}. For any $b,b^\prime\in A^2_{(2)}(X)$, there is equality
   \[  b\times b^\prime - b^\prime\times b=0\ \ \ \hbox{in}\ A^4(X\times X)\ .\]
   \end{proposition}
   
   \begin{proof} Let $\Psi^\prime\in A^{2}(B\times X)$ be inverse to $ \Gamma$ in the isomorphism (\ref{isomotch}), and let
    \[ \Psi:=   \pi^X_{2}\circ \Psi^\prime\circ \pi^B_{2,tr}\ \ \ \in A^{2}(B\times X)\ .\]
    It follows formally that $\Psi$ induces isomorphisms
    \[ \Psi_\ast\colon\ \ \ A^2_{(2)}(B)\ \xrightarrow{\cong}\ A^2_{(2)}(X)\ .\]
    There is a commutative diagram
    \[ \begin{array}
       [c]{ccc}
       A^2_{(2)}(B)\otimes A^2_{(2)}(B) & \xrightarrow{\times} &A^4_{(4)}(B\times B)\\
      \ \ \ \ \ \ \ \ \ \ \ \ \ \  \downarrow{\scriptstyle (\Psi_\ast,\Psi_\ast)} && \ \ \ \ \ \ \ \  \ \ \ \ \downarrow{\scriptstyle (\Psi\times \Psi)_\ast}\\
       A^2_{(2)}(X)\otimes A^2_{(2)}(X) & \xrightarrow{\times} &A^4_{}(X\times X)\\
       \end{array}\]
       and (as we have just seen) the left--hand vertical arrow is an isomorphism. To prove proposition \ref{intermed}, it thus suffices to prove the corresponding statement for the abelian surface $B$; this is taken care of by a result of Voisin's:
       
     \begin{theorem}[Voisin \cite{Vo}]\label{thabvar} Let $B$ be an abelian variety of dimension $g$. For any $c,c^\prime\in A^g_{(g)}(B)$, there is equality
     \[  c\times c^\prime -(-1)^g\ c^\prime\times c=0\ \ \ \hbox{in}\ A^{2g}_{(2g)}(B\times B)\ .\]
     \end{theorem}
       
    (This is \cite[Example 4.40]{Vo}. The notation $A^\ast_{(\ast)}$ refers to Beauville's splitting of the Chow ring of an abelian variety \cite{Beau}.)
    
    To go from proposition \ref{intermed} to theorem \ref{main}, we apply theorem \ref{chowringfano}. Let us first suppose $j=2$, i.e. $a,a^\prime$ are zero--cycles in $A^4_{(2)}(X)$. According to theorem \ref{chowringfano}(\rom1), we can express $a$ and $a^\prime$ as
     \[   a= b\cdot \ell\ \ \ , a^\prime=b^\prime\cdot\ell\ ,\]
     for some (unique) $b,b^\prime\in A^2_{(2)}(X)$. It follows that
     \[ \begin{split} a\times a^\prime - a^\prime\times a &= (b\cdot\ell)\times (b^\prime\cdot\ell) - (b^\prime\cdot\ell)\times (b\cdot\ell)\\
                                                                                 &= (b\times b^\prime - b^\prime\times b)\cdot (\ell\times\ell)\\
                                                                                 &=0\ \ \  \hbox{in}  \ A^8(X\times X)\ ,\\
                                                                          \end{split}\]
                                                              where the last line follows from proposition \ref{intermed}. 
                                                              
          Next, let us suppose $j=4$, i.e. $a,a^\prime$ are zero--cycles in 
                                                              $A^4_{(4)}(X)$. According to theorem \ref{chowringfano}(\rom2), we can express $a$ and $a^\prime$ as
   \[ a=b_1\cdot b_2\ \ \ ,\ a^\prime=b_1^\prime\cdot b_2^\prime\ ,\]
   for some $b_1,b_2,b_1^\prime,b_2^\prime\in A^2_{(2)}(X)$. It follows that
   \[ \begin{split} a\times a^\prime  &= (b_1\cdot b_2)\times (b_1^\prime\cdot b_2^\prime)     \\
                                                                                 &= (b_1\times b_1^\prime)\cdot   (b_2\times b_2^\prime) \\
                                                                                 &= (b_1^\prime\times b_1)\cdot   (b_2^\prime\times b_2) \\
                                                                                 &= (b_1^\prime\cdot b_2^\prime) \times (b_1\cdot b_2)\\                                                                                 
                                                                                 &=   a^\prime\times a   \ \ \  \hbox{in}  \ A^8(X\times X)\ ,\\
                                                                       \end{split}\]
                                                                       where in the middle we have used proposition \ref{intermed}.
                                         This settles the case $j=4$. The remaining case $j=0$ is trivially true since $A^4_{(0)}(X)\cong\QQ$.          
    \end{proof}

 \begin{proposition}\label{fanofindim} Let $Y\subset\PP^5(\C)$ be a cubic as in theorem \ref{main}, and let $X=F(Y)$ be the Fano variety of lines in $Y$. Then $Y$ and $X$ have finite--dimensional motive.
  \end{proposition}  
    
 \begin{proof} To establish finite--dimensionality of $Y$ is an easy exercice in using what is commonly known as the ``Shioda inductive structure'' \cite{Shi}, \cite{KS}. Indeed, applying 
 \cite[Remark 1.10]{KS}, we find there exists a dominant rational map
   \[ \phi\colon\ \ \ Y_1\times Y_2\ \dashrightarrow\ Y\ ,\]
   where $Y_1, Y_2\subset\PP^3(\C)$ are smooth cubic surfaces defined as
     \[ \begin{split} &f(X_0,X_1,X_2)+V^3=0\ ,\\
                          &g(X_0,X_1,X_2)+W^3=0\ .\\
                       \end{split}\]
           The indeterminacy locus of $\phi$ is resolved by blowing up the locus $C_1\times C_2\subset Y_1\times Y_2$, where $C_1, C_2$ are cubic curves. As $Y_1,Y_2,C_1,C_2$ have finite--dimensional motive, the blow-up also has finite--dimensional motive. Since this blow--up dominates $Y$, the cubic $Y$ has finite--dimensional motive. 
           
       The main result of \cite{fano} now implies that the Fano variety $X=F(Y)$ also has finite--dimensional motive.            
        \end{proof}   
    
  \end{proof}

 \begin{corollary}\label{cor} Let $X$ be as in theorem \ref{main}. The Hodge sub--structure
     \[ \wedge^2 H^4(X)\ \subset\ H^8(X\times X) \]
     is supported on a divisor.
   \end{corollary}

 \begin{proof} As noted in \cite{V9}, this follows Bloch--Srinivas style from the truth of conjecture \ref{conjvois}. The idea is as follows. Let 
   \[\iota\colon X\times X\to X\times X\]
   denote the involution switching the two factors. The correspondence
   \[ \wedge^2 \pi_4^X := \bigl( \Delta_{X\times X} -\Gamma_\iota\bigr)\circ (\pi^X_4\times \pi^X_4)\ \ \ \in A^8(X^4) \]
   acts on cohomology as a projector on $\wedge^2 H^4(X)\ \subset\ H^8(X\times X)$. 
   On the other hand, the $j=4$ case of theorem \ref{main} implies that
     \[  (\wedge^2 \pi^X_4)_\ast A^8(X\times X) =0\ .\]
   Using the Bloch--Srinivas argument \cite{BS}, this implies the correspondence $\wedge^2 \pi^X_4$ is rationally equivalent to a cycle supported on $X\times X\times D$ where $D\subset X\times X$ is a divisor. Returning to the action of $\wedge^2 \pi^X_4$ on cohomology, this implies
   \[    \wedge^2 H^4(X)=(   \wedge^2 \pi^X_4)_\ast H^8(X\times X) \]
   is supported on $D$.    
 \end{proof}
 
 \begin{remark}
 \normalfont Using the work of Rie\ss\, \cite{Rie}, it is possible to extend theorem \ref{main} to all hyperk\"ahler varieties birational to the Fano variety $X$ of theorem \ref{main}.
 \end{remark}
 
 \begin{remark}\normalfont The family of Fano varieties of theorem \ref{main} has been much studied, starting with Namikawa's work \cite{Nam}. It also appears in \cite{Kawa}, and in
  \cite[Proposition 12]{Beau2} (where it is proven that for appropriate choices of $f$ and $g$, the Fano variety $X$ is ``$N^2$--maximal'' in the sense that
   \[ \dim_{\QQ} N^2 H^4(X) =\dim_\C H^{2,2}(X,\C)\ .)\] 
  \end{remark} 
 
 \begin{remark}\normalfont As the expert reader will have noticed, the crux of the proof of theorem \ref{main} is the existence of a symplectic automorphism of $X$ for which the fixed locus is an abelian surface. As such, it is natural to ask whether there are other families of Fano varieties of cubic fourfolds with this property. 
 
 However, if one adds the condition that the automorphism is of primary order and polarized, the Fano varieties as in theorem \ref{main} are the only Fano varieties of cubic fourfolds with this property (this follows from the classification given in \cite{LFu3}; the case studied in theorem \ref{main} appears as ``family IV-2'' in \cite[Theorem 0.1]{LFu3}, and there are no other instances where the fixed locus is an abelian surface).
 
 \end{remark}

\vskip1cm
\begin{nonumberingt} Thanks to all participants of the Strasbourg 2014/2015 ``groupe de travail'' based on the monograph \cite{Vo} for a very stimulating atmosphere.
Many thanks to Kai and Len for always respecting the rule not to bring toys into my office.
\end{nonumberingt}

\end{document}